OpenStudy (littlediva12459):

A bag contains 2 white marbles and 7 purple marbles. Two marbles are drawn at random. One marble is drawn and not replaced. Then a second marble is drawn. What is the probability of selecting a purple marble and then a white marble? @Aveline

OpenStudy (littlediva12459):

A. 7/36 B. 1/7 C. 1/14 D. 2/9

OpenStudy (aveline):

The chance of selecting a purple marble is 7/9. After we remove one marble, the chance of getting a white marble is 2/8. What is:\[\frac{ 2 }{ 8 } \times \frac{ 7 }{ 9 }\]?

OpenStudy (littlediva12459):

it would be 7/36 soA.

OpenStudy (littlediva12459):

thanks for the help